OpenStudy (surryyy):

We toss three coins. What is the probability of the event "3 heads appear"? Answer 1/4 3/8 1/8 1/2

OpenStudy (gorv):

well watzz ur say ..how many time 3 head can appear??

OpenStudy (surryyy):

the answer is 1/8?

OpenStudy (gorv):

yeah thatzz right ,,but how u got ??

OpenStudy (gorv):

@surryyy

OpenStudy (gorv):

tell me i know how bt just wanted to chck that u did right

OpenStudy (mayankdevnani):

When we toss 1 coin,the maximum probabilty we can get is 1 i.e head or tail. but there are total 2 outcomes. so probability :- \[\large \bf probability=\frac{no.of~favourable~conditions}{No.of~outcomes}=\frac{1}{2}\] but we tossed three times,so \[\large \bf probability=\frac{1}{2} \times \frac{1}{2} \times \frac{1}{2} =\frac{1}{8}=\color{red}{Answer}\]

OpenStudy (mayankdevnani):

hope you understand. @surryyy

OpenStudy (surryyy):

yea i did the same thing. but i wasnt sure if i did it right or not.

OpenStudy (mayankdevnani):

you are correct :) i mean your answer is correct

OpenStudy (mayankdevnani):

good job :) @surryyy

OpenStudy (gorv):

u r right bud...be confident :)

OpenStudy (mayankdevnani):

yeah... ^^^^

OpenStudy (surryyy):

Thnx.

OpenStudy (mayankdevnani):

welcome :)